### v2

# High-precision direct decay energy measurements of the electron-capture decay of 97Tc

Abstract
A direct measurement of the ground-state-to-ground-state electron-capture decay () value of 97Tc has been conducted employing the high resolving power phase-imaging ion-cyclotron-resonance technique with the double Penning trap mass spectrometer JYFLTRAP. The resulting value for 97Tc is 324.82(21) keV, exhibiting a precision approximately 19 times higher than the value adopted in the newest Atomic Mass Evaluation (AME2020) and differing by 1.2. Furthermore, by combining this refined value with nuclear energy-level data for the decay-daughter 97Mo, a potential ultra-low Q-value transition 97Tc (9/2+, ground state) 97Mo∗ (320(1) keV), was detected. The ground-state-to-excited-state electron-capture decay value () of this transition was determined to be 4.8(10) keV, confirming it to be energetically allowed with a confidence level of exceeding 4. The captures of electrons occupying the L and higher shells for this transition are energetically allowed, giving a value of 2.0(10) keV for the closest distance of to the allowed binding energy of the L1 shell. To predict partial half-lives and energy-release distributions for this transition, the atomic self-consistent many-electron Dirac–Hartree–Fock–Slater method and the nuclear shell model have been employed. Dominant correction terms such as exchange and overlap corrections, as well as shake-up and shake-off effects, were included in the final results. Moreover, in the case of a possible allowed transition, the normalized distribution of released energy in the electron-capture decay of 97Tc to the excited 320-keV state of 97Mo, is compared with that of 163Ho, which is being used for electron-neutrino-mass determination. A pseudo-experiment technique was introduced to calculate error propagation in half-life and the 68% confidence interval for normalized energy distributions.
I Introduction
Neutrinos are electrically neutral fermions that interact only via the weak force and gravity, and they have an incredibly tiny rest mass compared to other elementary particles. Neutrinos do not participate in electromagnetic or strong interactions, thus allowing them to pass through matter unimpeded. Up to now, three neutrino flavors have been found: electron neutrino (), muon neutrino (), and tau neutrino (). For each neutrino, there also exists a corresponding antineutrino, which also has spin of 1/2 and no electric charge Petcov (2013). Although neutrinos were initially thought to be massless as assumed by the standard model (SM), the three neutrino mass eigenstates do not show a one-to-one correspondence to the three flavors, each flavor being a quantum superposition of all three mass states Lesgourgues and Pastor (2006). Neutrinos oscillate between flavors, making their detection even more intriguing Fukuda et al. (1998); SNO Collaboration (2002); Gerbino and Lattanzi (2018). Neutrino-oscillation experiments reveal that at least two neutrino-mass eigenstates have a non-zero rest mass, leading to a mass difference between the three mass eigenstates. Laboratory experiments and cosmological observations have provided experimental data on differences in the squares of the neutrino masses, an upper limit on their sum, and an upper limit on the mass of the electron (anti)neutrino. Among these experiments, the most direct and model-independent way to pin down the exact mass value of the electron antineutrino (m) is to measure the electron energy spectrum of single decay, e.g., the Karlsruhe Tritium Neutrino (KATRIN) experiment Drexlin et al. (2013); Aker et al. (2019, 2022), which has recently pushed the absolute electron antineutrino mass scale boundary to 0.45 eV/c2 at the 90% confidence level (CL) KATRIN Collaboration et al. (2025). KATRIN will finally set the m upper limit with a sensitivity of 0.2 eV/c2 at 90% CL. Other single--decay experiments, Project 8 and PTOLEMY project, using cyclotron radiation emission spectroscopy and the development of atomic tritium sources, will eventually allow a sensitivity of tritium-based neutrino-mass experiments to m beyond the design sensitivity of KATRIN Esfahani et al. (2017); Betti et al. (2019) down to 0.04 eV/c2. An alternative method, employing electron capture in 163Ho (ECHo), an experiment to calorimetrically measure the electron-capture (EC) deexcitation spectrum Gastaldo et al. (2014, 2017); Velte et al. (2019), has set a current limit of 150 eV/c2 for the electron neutrino mass Velte et al. (2019).
In single--decay experiments aimed at determining the mass of the electron antineutrino, a value of the decay as small as possible is sought. A smaller value leads to a higher fraction of effective decays within a given energy interval near the endpoint region. Currently, the running direct (anti)neutrino-mass-determination experiments rely exclusively on ground-state-to-ground-state (gs-to-gs) decay cases, such as 3H ( decay) and 163Ho (EC). However, ongoing intensive searches are actively exploring isotopes having low -value ground-state-to-excited-state (gs-to-es) /EC decays Cattadori et al. (2007); Haaranen and Suhonen (2013); Suhonen (2014); Ejiri et al. (2019) using high-precision Penning-trap mass spectrometry (PTMS) including JYFLTRAP, LEBIT, CPT, ISOTRAP, and SHIPTRAP Penning traps Sandler et al. (2019); Karthein et al. (2019); De Roubin et al. (2020); Ge et al. (2021a, b, 2022a); Eronen et al. (2022); Ge et al. (2022a, b); Ramalho et al. (2022); Gamage et al. (2022); Keblbeck et al. (2023); Ge et al. (2023); Valverde et al. (2024); Ge et al. (2024a, b). PTMS is the most precise and accurate technique for determining atomic masses and values. It remains the only direct method capable of measuring decay values to sub-keV precision or better, allowing verification of potential ultra-low (¡ 1 keV) -value transitions.
For the nuclide 97Tc, there exists a potential low -value gs-to-es EC transition, 97Tc (9/2+, ground state) 97Mo∗ (320(1) keV NND (2021), excited state), that could be employed for neutrino-mass detection. This speculated allowed transition, having a gs-to-es -value of 0.0(41) keV deduced from the existing experimental data NND (2021); Wang et al. (2021), makes itself a candidate for an ultra-low -value transition. However, the value of the decay to the excited state in the daughter nucleus is known with only 4.1 keV precision. This uncertainty arises from the combined uncertainties in the gs-to-gs EC value (4 keV) adopted from Wang et al. (2021) and in the excitation energy (1 keV) from NND (2021). As a result, a definitive conclusion regarding its energetic feasibility cannot be drawn solely from the current experimental information. To further determine whether this specific decay is energetically feasible and whether it falls into the ultra-low category, experimental measurements of the gs-to-gs decay value and the excitation energy of the state in the daughter nucleus are required.
Here, we present the first direct measurement of the gs-to-gs EC value for 97Tc using the JYFLTRAP Penning-trap mass spectrometer. The precise value obtained, along with current nuclear energy-level information for excited states of 97Tc, is employed to refine the gs-to-es values. The possible ultra-low -value transition in 97Tc is verified with the newly determined experimental data. We confirm the transition to be energetically allowed. To explore the viability of this transition for neutrino-mass detection, we utilize two computational methods: the atomic self-consistent many-electron Dirac–Hartree–Fock–Slater method and the nuclear shell model. These approaches enable us to predict the partial half-lives and energy-release distributions for the relevant EC-decay transition.
II Experimental method
The experiment was carried out at the Ion Guide Isotope Separator On-Line facility (IGISOL) at the University of Jyväskylä, Finland Moore et al. (2013); Kolhinen et al. (2013). Figure 1 gives a schematic view of the experimental setup.
97Tc ions were generated in fusion-evaporation reactions between a natural molybdenum target foil irradiated with a few A of a proton beam of energy 45 MeV produced by the K-130 cyclotron. The secondary ions were stopped in a small volume gas cell filled with helium at a pressure of around 100 mbar as shown in Fig. 1. There, the ions end up mainly singly charged from charge-exchange reactions. The predominantly singly charged ions are extracted from the gas cell by helium gas flow and are guided with electric fields facilitated by differential pumping and a sextupole ion guide (SPIG) Karvonen et al. (2008). Subsequently, the ions were accelerated using a 30 kV electric potential and transported to the 55∘ dipole magnet for isobaric mass separation with a typical mass resolving power of approximately 500. The isobarically separated ions with mass-to-charge ratio = 97 (including reaction products such as 97Nb+, 97mTc+, 97Tc+, and 97Mo+) were then directed to a radiofrequency-quadrupole cooler-buncher (RFQ-CB) Nieminen et al. (2001), where they underwent accumulation, cooling, and bunching. To produce the decay-daughter ion 97Mo+ samples, an offline glow-discharge ion source was used Vilén et al. (2020). Additionally, a 90∘ electrostatic bender, as presented in Fig. 1, selectively guided ions either from the online target station or the offline ion source for downstream transmission to the RFQ-CB.
The cooled and bunched ions from the RFQ-CB, were further directed to the JYFLTRAP Eronen et al. (2012), which consists of two cylindrical Penning traps within a 7-T superconducting solenoid, for the final mass and -value measurements. The first trap, filled with helium buffer gas, serves as a purification trap and is employed for isobaric purification using the sideband buffer gas cooling technique Savard et al. (1991) with a resolving power of approximately . To select the 97Tc+ ion samples, this method could separate the aforementioned ions of 97Nb+, but lacked sufficient mass resolving power to eliminate 97Mo+ and 97mTc+ ions.
Mono-isotopic samples of 97Tc+ were therefore prepared using the coupling of the dipolar excitation with Ramsey’s method of time-separated oscillatory fields Eronen et al. (2008) and the phase-imaging ion-cyclotron-resonance (PI-ICR) technique Nesterenko et al. (2021, 2018), as described in Ge et al. (2023). A Ramsey-type dipole excitation frequency scan in the second trap, employing a 2 ms (On) - 98 ms (Off) - 2 ms (On) excitation pattern, was used. The ions of interest were selectively chosen by filtering through positional gates using PI-ICR identification with a 436 ms phase accumulation time.
The PI-ICR method Nesterenko et al. (2018, 2021) is used to measure the value via the actual measurements of the cyclotron frequency,
where is the mass of the ion, the charge and the magnetic field of the trap. The PI-ICR technique is around 25 times faster than the conventional time-of-flight ion-cyclotron-resonance (TOF-ICR) method used to achieve the same precision with a similar experimental condition Nesterenko et al. (2018, 2021); Eliseev et al. (2014, 2013). The scheme of the PI-ICR technique used at JYFLTRAP Nesterenko et al. (2018) relies on the direct measurements of the magnetron motion and the reduced cyclotron motion at the same time while projecting the radial ion motion onto a position-sensitive MCP detector. The angle between the two phase images of the projected radial motions, obtained from two timing patterns for determining the cyclotron frequency , is defined with respect to the center spot: . Specifically, and correspond to the polar angles of the magnetron and reduced cyclotron motion phases. The cyclotron frequency is determined from the measurements during the phase accumulation time :
where represents the full number of revolutions of the measured ions. To unambiguously assign in Eq. (2), different accumulation times were used for 97Tc+. The actual measurements employed an accumulation time of 436 ms to determine the final cyclotron frequency for both 97Tc+ and 97Mo+ ions, ensuring that any leaked isobaric contaminant would not overlap with the ions of interest.
To minimize any shifts in the ratio of the 97Tc+-97Mo+ pair, the positions of the phase spots for the magnetron and reduced cyclotron motions were carefully chosen to maintain the angle within a few degrees. This was done to account for the conversion of the reduced cyclotron motion to the magnetron motion and the potential distortion of the ion-motion projection onto the detector, ensuring that the shift remained well below 10-10 Eliseev et al. (2014). Figure 2 shows a representative measurement with ”reduced cyclotron” and ”magnetron” phase spots relative to the center spot. During the measurements, the excitation delay of was systematically scanned over one magnetron period, while the extraction delay varied over one cyclotron period, accounting for any residual motion that might have affected the spots. The interleaved measurements of for 97Tc+-97Mo+ ions had a total data accumulation time of approximately 4.9 hours.
The gs-to-gs EC value (), the mass difference of the decay pair, can be derived from their measured cyclotron frequency ratio (=):
where and are the masses of the parent and daughter atoms, and the mass of an electron. The electron binding energies of the parent and daughter atoms, denoted as and , are neglected due to their small values (on the order of a few eVs Kramida et al. (2020)), and is close to 1. Given that both the parent and daughter ions share the same value and their relative mass difference , the mass-dependent error becomes insignificant compared to the statistical uncertainty achieved in the measurements. Additionally, the contribution of uncertainty to the value arising from the mass uncertainty of the daughter ion chosen as a reference, 0.16 keV/c2 for 97Mo Wang et al. (2021), can be safely disregarded.
| 97Tc (AME2020) | 320.0(40) | -87224.0(40) | ||
| 97Tc (This Work) | 1.00 000 359 84(23) | 324.82(21) | -87219.88(26) |
| Final state |
| 97Mo () | 320.0(40) | 324.82(21) | 0.0(41) | 4.8(10) | 320.0(10) | 4.8 | 2.0(10) | 2.2(10) | 4.3(10) | ||||||||||||||||
| 97Mo (5/2+, ground state) | 320.0(40) | 324.82(21) |
III Results and discussion
From the measurements of the cyclotron frequency ratio , was calculated via Eq. 3. As shown in Fig. 3, 20 PI-ICR measurement data sets were collected for 97Tc+-97Mo+. In less than 5 minutes, a full scanning measurement (one cycle) of the magnetron phase, reduced cyclotron phase and center spot in sequence was completed for each ion species of the decay pair 97Tc+-97Mo+. In the analysis, we determined the position of each spot using the maximum likelihood method. After summing a few cycles to obtain reasonable counts for fitting, phase angles were calculated based on the determined positions of the phase spots as demonstrated in Fig. 2. These phase angles allowed us to deduce the cyclotron frequencies of each ion species. By linearly interpolating the of the daughter ion 97Mo+ (reference) to the time of the measurement of the parent ion 97Tc+ (ion of interest), the cyclotron frequency ratio was derived. To minimize potential cyclotron frequency shifts resulting from ion-ion interactions Kellerbauer et al. (2003); Roux et al. (2013), a median ion rate was constrained to 1-2 ions per bunch during the measurements. Bunches with fewer than five detected ions per bunch were considered during the data analysis. In the analysis, the count-rate related frequency shifts were not observed. The temporal fluctuation of the magnetic field /min Nesterenko et al. (2021) is taken in account in the final results. Here, represents the time interval between consecutive reference measurements. The contribution of magnetic field fluctuations to the final frequency ratio uncertainty was less than 10-10, as the parent-daughter measurements were interleaved with intervals less than 10 minutes. Additionally, the frequency shifts in the PI-ICR measurement caused by ion image distortions were ignored, as they were well below the statistical uncertainty. Furthermore, the fact that decay pair ions 97Tc+-97Mo+ are mass doublets, cancels out many systematic uncertainties in the cyclotron frequency ratio. The weighted mean ratio of all single frequency ratios was calculated, considering the inner and outer errors to deduce the Birge ratio Birge (1932). The maximum of the inner and outer errors served as the weight for computing . Results of our analysis including all data sets, in comparison to literature values, are presented in Fig. 3 and Table 1. The final daughter-to-parent frequency ratio with its uncertainty is determined to be 1.000 003 598 4(23), resulting in a gs-to-gs value of 324.82(21) keV.
The gs-to-gs value of 324.82(21) keV obtained from this work, which represents the first direct measurement, is approximately 19 times more precise than the value deduced from the evaluated masses in AME2020 Huang et al. (2021); Wang et al. (2021). The measured value has a deviation of 4.8(40) keV from the AME2020 value and is 1 larger.
In the case of EC, the closer is the value of the decay to one of the ionization energies of the captured electrons, the larger the resonance enhancement of the rate near the end-point, where the effects of a non-vanishing neutrino mass are relevant. Which orbital electrons take part in the EC process and the absolute values of the decays are crucial for modelling the spectrum shape near the endpoint. The current precision deduced from literature NND (2021); Huang et al. (2021); Wang et al. (2021) does not allow a definitive conclusion on whether any other electron orbitals allow EC to occur for the transition 97Tc (9/2+) 97Mo∗ (320.0 keV). In this work, the high-precision -value measurement allows an unambiguous characterization of all the possible lines in the EC spectrum at a significance level of at least 2 for the transition. This makes the modeling of its shape possible. The high-precision EC energy measured from this work, together with the nuclear energy level data from Ref. NND (2021) of the excited state of 97Mo as tabulated in Table 2, was used to determine the gs-to-es value () of the candidate state. The newly determined confirms that the decay of the ground state of 97Tc to the excited state of interest is energetically allowed with a confidence level of more than 4, as shown in Table 2. A comparison of the values of the gs-to-es EC transition from this work to the value derived from AME2020 is shown in Fig. 4. As tabulated in Table 2, gives the distance of the value to the binding energy X-R (2021) of the electrons in the allowed daughter atomic shells ( = L1, L2, and other electrons from s levels and p1/2 levels from the third and higher shells). For the state with the excitation energy of 320.0(10) keV, the captures of electrons occupying the L1 and higher shells for the transition 97Tc (9/2+) 97Mo∗ are energetically allowed as indicated in Fig. 4. The transition 97Tc (9/2+) 97Mo∗ (320.0 keV), giving the values of 2.0(10) keV and 2.2(10) keV for the distance of to the allowed binding energies = 2.866 keV and = 2.625 keV X-R (2021), is potentially of the allowed Fermi plus Gamow-Teller type. This transition is of high interest for future neutrino mass determination due to a possible high branching ratio. To confirm whether the emitted neutrino energy 2.0(10) keV is ultra-low, further high-precision measurements of the excitation energy of the state are required. Moreover, the parity of this state needs to be determined to verify the decay type of the transition to this state.
IV Theoretical predictions
Two calculation methods are employed to predict the distribution of energy released in the decay, the transition half-life and their error spread, namely the atomic many-electron Dirac–Hartree–Fock–Slater (DHFS) self-consistent method and the Nuclear Shell Model (NSM) many-nucleon framework. For the atomic-structure calculations we made use of the RADIAL subroutine package Salvat and Fernández-Varea (2019), which also contains the DHFS.F code, while for the NSM we employed the NuShellXMSU code Brown and Rae (2014).
In our calculation we consider the initial atom in its ground state. The electronic structure of the final atom is constructed starting from the configuration of the initial atom with an electron hole in the shell from where it was captured. When the nucleus captures the electron, the other electrons (the spectator electrons) undergo processes that impact the decay rate. A couple of important corrections related to these processes are considered Ge et al. (2024b): exchange and overlap corrections, shake-up and shake-off effects.
The atomic shells are denoted using the main quantum number , and the relativistic quantum number as . In order to compute the released energy distribution of an allowed EC event we sum all the contributions from the atomic shells with the relativistic quantum number as:
| LCS | Decay | Total half-life | ||||||||
| type | ||||||||||
| L1 | allowed | |||||||||
| L2 | allowed | |||||||||
| Total half-life | ||||||||||
| L1 | 2nd UF | |||||||||
| L2 | 2nd UF | |||||||||
| Total half-life | ||||||||||
| L1 | 4th NUF | |||||||||
| L2 | 4th NUF |
Here we employ two corrections, namely the exchange and overlap correction represented by the factor and the shake-up and shake-off corrections introduced as the factor Mougeot (2018). The Coulomb amplitude is calculated for the initial atom and is represented by . The Cabibbo angle and the Fermi constant are present in . The relative occupancy of the shell is denoted as . The energy and momentum of the emitted neutrino are represented as and , with being the neutrino mass. A recent improvement Sevestrean et al. (2023) in the computation of the relaxation energy following the capture is the usage of total binding energy as , where indicates the energy of the final atom in its ground state, while represents the energy of the excited state of the final atom with a hole in the shell. The represents the intrinsic line-widths of Breit–Wigner resonances centered at and they were taken from CAMPBELL and PAPP (2001).
The spin of the 320-keV excited state in 97Mo is uncertain, ranging within a range of . This means that the decay transition from the ground state of 97Tc to the 320-keV state could be allowed ( and ), second-forbidden non-unique (2nd NUF) in the case of , second-forbidden unique (2nd UF) in the case of and fourth-forbidden non-unique (4th NUF) in the case of . In the case of an allowed transition, the nuclear structure part is represented by the factor
with being the weak vector coupling, and and being the weak axial coupling and the angular momentum of the initial nucleus. The Fermi nuclear matrix element is denoted by and the Gamow-Teller nuclear matrix element is denoted as .
To compute the decay probability we employ the formalism described in Bambynek et al. (1977):
where . For allowed transition we used the simplified approach for the factor in Eq. 5, while for the unique and non-unique forbidden transitions, is defined as:
The and factors are defined in Bambynek et al. (1977).
In Table 3, we present the predicted half-lives and associated uncertainties for the decay of 97Tc to the excited state of 97Mo, corresponding to the experimental energy keV, across all relevant atomic shells. Due to the uncertainty of the angular momentum of the final state, we considered multiple possible values. By computing the energy levels using the jj45pnb interaction, we obtained a few possible values for J in the vicinity of 320 keV, namely 7/2, 3/2, and 1/2. Thus, we have the possible decay types allowed, 2nd unique forbidden, and 4th non-unique forbidden. The matrix elements needed for the half-life were computed with the same jj45pnb interaction. We also used the glekpn interaction for the allowed transition; the computed half-life was an order of magnitude smaller. Comparing the half-life for the excited state transition with the total half-life measured, years NND (2021), one observes that for the allowed transition, the computed half-life is in tension with the measured one. As such, this is most likely due to the assumed angular momentum of the excited final state, which thus should be verified by new spectroscopic measurements. Considering the uncertainty in the value, the capture from the L1 shell is uncertain but we retain this possibility in our computations. However, the capture from the L2 shell is certainly possible. As can be seen, there is a significant difference between the upper and lower error because the -value spans a considerable interval of values, and it appears at various powers in the half-life formula.
In Fig. 5 we present the normalized distributions of released energy within a shaded-in-blue 68 confidence band between the blue curves assuming an allowed transition. Considering a Gaussian distribution for the Q-value, 97.3 of cases permit the capture of an electron from the L1 shell (bottom plot). In 1.2 the L1 channel is energetically impossible but the capture from the L2 sub-shell is possible (top plot). The other 1.5 of cases were not presented here, due to their decreasing probability of occurrence. The black line represents the median values for each energy point in the graph. The interval was obtained as described in section IV.1. This analysis shows the high dependence of the normalized distribution on the value of , especially in the interesting region at the end of the spectrum. The greatest impact on the error of comes from the uncertainty in the energy of the level of interest in the final nucleus. Thus a better assessment of the energy of this level is necessary in order to have a more well-defined distribution of energy at .
The current measurement of the value of with the associated error does not give a definite picture of the capture process. In order to investigate the implications of the associated uncertainties in the case of an allowed transition, we present in Fig. 6 a comparison between the normalized distribution of the released energy as a function of the energy, , for the decay of 97Tc to the excited state of 97Mo with keV, within 1 range of the experimental central value, and 163Ho to the ground state of 163Dy with keV. As can be seen in the enlarged scale in the inset of Fig. 6, the neutrino-mass-sensitive region provides a higher EC event rate for 97Tc for this Q value, making it comparable with the rate for 163Ho. Also, this inset shows exactly the impact of the neutrino mass on the shape of the spectrum and how it could be highlighted. A strong 97Tc source can be produced via the reaction with 19 MeV protons on an enriched 97Mo target (cross-section 206 mbarn from theoretical calculation with TALYS Koning, Arjan et al. (2023)). With a 10 mA beam and 1 month irradiation, we estimate producing atoms, sufficient for a neutrino mass experiment requiring sub-keV sensitivity over one year if the allowed transition can be confirmed. The long half-life of about 4 My for 97Tc renders it a good candidate for future long-term neutrino-mass measurements if the considered transition is of the allowed type. In case the transition is of forbidden type, it is not relevant for neutrino-mass measurements as witnessed by the very long half-lives of the transition in Table 3.
IV.1 Statistical approach
In order to calculate the propagation of errors in the half-life and the 68% band in the normalized distributions of released energy as shown in Fig. 5, we employed the pseudo-experiment technique. This method involves sampling of parameters that potentially fluctuate from probability distribution functions (PDFs). For each set of parameters, the half-life times and the normalized distribution of released energy are computed, thus obtaining their PDFs. The central values are obtained as medians of the PDFs. For the two-sided bounds, we took the element at and of the CPDFs. The uncertainties are reported as the difference between each bound and the median value.
We took into account and as sources of uncertainty. Assuming a Gaussian distribution of , in about cases the capture from the L1 sub-shell will not be energetically possible. Thus, we split the sampled values into groups and considered two in the current study. For the first group, the capture of the L1 electron is possible (L1 allowed), while for the second, the L1 channel is energetically forbidden, and the L2 channel is energetically possible (L1 forbidden). For each group, we plotted the median and the 1 band of the normalized distribution of released energy. For the distribution of we have used a uniform distribution with values between and . The error propagation based on was taken into account only for the allowed transition. We consider a number of pseudo-experiments. In particular, for the normalized distributions of released energy, we constructed a uniform grid of points for . In each point, the procedure described above was applied.
V Conclusion and outlook
A direct high-precision gs-to-gs EC-decay -value measurement of 97Tc (9/2+)97Mo (5/2+) was performed using the PI-ICR technique at the JYFLTRAP Penning-trap mass spectrometer. A value of 324.82(21) keV was obtained and the precision was improved by a factor of around 19 compared to literature. The measurement also improved the mass excess of 97Tc , -87219.88(26) keV/c2, by a factor of 15 compared to the previously measured value. With the refined gs-to-gs value, the candidate transition of 97Tc (9/2+)97Mo∗ (320.0 keV) was validated to be energetically allowed. This also allows an unambiguous characterization of all the possible lines in the EC spectrum of this transition at a significance level of at least 2. The transition to the 320.0-keV state has a small distance of 2.0(1.0) keV for the gs-to-es to the binding energy of the electrons in the allowed daughter atomic shell L1. For further confirmation of whether this lowest emitted neutrino energy is ultra-low, the excitation energy needs to be determined with higher precision and accuracy. Also, to define the nature of the transition, allowed or forbidden, a spectroscopic measurement of the angular momentum of the final state should be performed. A highly charged ion (HCI) Penning trap will refine the gs-to-gs to a few eV Schweiger et al. (2024), and cryogenic detectors, e.g., Transition Edge Sensors (TES) or Metallic Magnetic Calorimeters (MMCs) at low temperatures, will measure the 320.0 keV excitation energy to around 10s eV, reducing uncertainty to 10s eV. Phase-space factor analysis and gamma coincidence further constrain rates, ensuring viability for sub-eV sensitivity, as with 159Dy Gastaldo (2022). The atomic self-consistent many-electron Dirac–Hartree–Fock–Slater method and the nuclear shell model were utilized to predict the partial decay half-lives and energy distributions of gs-to-es EC transitions in 97Tc with low values. To calculate error propagation in half-life and the 68% confidence interval for normalized energy distributions, a pseudo-experiment technique was introduced. Multiple corrections, such as exchange, overlap, shake-up, and shake-off effects, were accounted for in these predictions for modeling the spectrum shape near the endpoint.
Acknowledgements.
We acknowledge the staff of the Accelerator Laboratory of University of Jyväskylä (JYFL-ACCLAB) for providing stable online beam. We thank the support by the Academy of Finland with projects No. 306980, 312544, 275389, 284516, 295207, 314733, 315179, 327629, 320062, 354589, 345869 and 354968. The support by the EU Horizon 2020 research and innovation program under grant No. 771036 (ERC CoG MAIDEN) is acknowledged. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 861198–LISA–H2020-MSCA-ITN-2019. V.A.S., O.N., S.S., J.S., and J.K. acknowledge support from the NEPTUN project (PNRR-I8/C9-CF264, Contract No. 760100/23.05.2023 of the Romanian Ministry of Research, Innovation and Digitization). The work leading to this publication was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - AY 155/2-1. This project has received funding from the European Union’s Horizon Europe Research and Innovation programme under Grant Agreement No 101057511 (EURO-LABS).References
- Petcov (2013) S. T. Petcov, Adv. High Energy Phys. 2013, 1 (2013).
- Lesgourgues and Pastor (2006) J. Lesgourgues and S. Pastor, Physics Reports 429, 307 (2006).
- Fukuda et al. (1998) Y. Fukuda, T. Hayakawa, E. Ichihara, K. Inoue, K. Ishihara, H. Ishino, Y. Itow, T. Kajita, J. Kameda, S. Kasuga, K. Kobayashi, Y. Kobayashi, Y. Koshio, M. Miura, M. Nakahata, S. Nakayama, A. Okada, K. Okumura, N. Sakurai, M. Shiozawa, Y. Suzuki, Y. Takeuchi, Y. Totsuka, S. Yamada, M. Earl, A. Habig, E. Kearns, M. D. Messier, K. Scholberg, J. L. Stone, L. R. Sulak, C. W. Walter, M. Goldhaber, T. Barszczxak, D. Casper, W. Gajewski, P. G. Halverson, J. Hsu, W. R. Kropp, L. R. Price, F. Reines, M. Smy, H. W. Sobel, M. R. Vagins, K. S. Ganezer, W. E. Keig, R. W. Ellsworth, S. Tasaka, J. W. Flanagan, A. Kibayashi, J. G. Learned, S. Matsuno, V. J. Stenger, D. Takemori, T. Ishii, J. Kanzaki, T. Kobayashi, S. Mine, K. Nakamura, K. Nishikawa, Y. Oyama, A. Sakai, M. Sakuda, O. Sasaki, S. Echigo, M. Kohama, A. T. Suzuki, T. J. Haines, E. Blaufuss, B. K. Kim, R. Sanford, R. Svoboda, M. L. Chen, Z. Conner, J. A. Goodman, G. W. Sullivan, J. Hill, C. K. Jung, K. Martens, C. Mauger, C. Mc Grew, E. Sharkey, B. Viren, C. Yanagisawa, W. Doki, K. Miyano, H. Okazawa, C. Saji, M. Takahata, Y. Nagashima, M. Takita, T. Yamaguchi, M. Yoshida, S. B. Kim, M. Etoh, K. Fujita, A. Hasegawa, T. Hasegawa, S. Hatakeyama, T. Iwamoto, M. Koga, T. Maruyama, H. Ogawa, J. Shirai, A. Suzuki, F. Tsushima, M. Koshiba, M. Nemoto, K. Nishijima, T. Futagami, Y. Hayato, Y. Kanaya, K. Kaneyuki, Y. Watanabe, D. Kielczewska, R. A. Doyle, J. S. George, A. L. Stachyra, L. L. Wai, R. J. Wilkes, and K. K. Young, Physical Review Letters 81, 1562 (1998), arXiv:9807003 [hep-ex] .
- SNO Collaboration (2002) SNO Collaboration, Physical Review Letters 89, 1 (2002), arXiv:0204008 [nucl-ex] .
- Gerbino and Lattanzi (2018) M. Gerbino and M. Lattanzi, Frontiers in Physics 5 (2018), 10.3389/fphy.2017.00070.
- Drexlin et al. (2013) G. Drexlin, V. Hannen, S. Mertens, and C. Weinheimer, Advances in High Energy Physics 2013, 1 (2013), arXiv:1307.0101 .
- Aker et al. (2019) M. Aker, K. Altenmüller, M. Arenz, M. Babutzka, J. Barrett, S. Bauer, M. Beck, A. Beglarian, J. Behrens, T. Bergmann, U. Besserer, K. Blaum, F. Block, S. Bobien, K. Bokeloh, J. Bonn, B. Bornschein, L. Bornschein, H. Bouquet, T. Brunst, T. S. Caldwell, L. La Cascio, S. Chilingaryan, W. Choi, T. J. Corona, K. Debowski, M. Deffert, M. Descher, P. J. Doe, O. Dragoun, G. Drexlin, J. A. Dunmore, S. Dyba, F. Edzards, L. Eisenblätter, K. Eitel, E. Ellinger, R. Engel, S. Enomoto, M. Erhard, D. Eversheim, M. Fedkevych, A. Felden, S. Fischer, B. Flatt, J. A. Formaggio, F. M. Fränkle, G. B. Franklin, H. Frankrone, F. Friedel, D. Fuchs, A. Fulst, D. Furse, K. Gauda, H. Gemmeke, W. Gil, F. Glück, S. Görhardt, S. Groh, S. Grohmann, R. Grössle, R. Gumbsheimer, M. Ha Minh, M. Hackenjos, V. Hannen, F. Harms, J. Hartmann, N. Haußmann, F. Heizmann, K. Helbing, S. Hickford, D. Hilk, B. Hillen, D. Hillesheimer, D. Hinz, T. Höhn, B. Holzapfel, S. Holzmann, T. Houdy, M. A. Howe, A. Huber, T. M. James, A. Jansen, A. Kaboth, C. Karl, O. Kazachenko, J. Kellerer, N. Kernert, L. Kippenbrock, M. Kleesiek, M. Klein, C. Köhler, L. Köllenberger, A. Kopmann, M. Korzeczek, A. Kosmider, A. Kovalík, B. Krasch, M. Kraus, H. Krause, L. Kuckert, B. Kuffner, N. Kunka, T. Lasserre, T. L. Le, O. Lebeda, M. Leber, B. Lehnert, J. Letnev, F. Leven, S. Lichter, V. M. Lobashev, A. Lokhov, M. MacHatschek, E. Malcherek, K. Müller, M. Mark, A. Marsteller, E. L. Martin, C. Melzer, A. Menshikov, S. Mertens, L. I. Minter, S. Mirz, B. Monreal, P. I. Morales Guzmán, K. Müller, U. Naumann, W. Ndeke, H. Neumann, S. Niemes, M. Noe, N. S. Oblath, H. W. Ortjohann, A. Osipowicz, B. Ostrick, E. Otten, D. S. Parno, D. G. Phillips, P. Plischke, A. Pollithy, A. W. Poon, J. Pouryamout, M. Prall, F. Priester, M. Röllig, C. Röttele, P. C. Ranitzsch, O. Rest, R. Rinderspacher, R. G. Robertson, C. Rodenbeck, P. Rohr, C. Roll, S. Rupp, M. Ryšavý, R. Sack, A. Saenz, P. Schäfer, L. Schimpf, K. Schlösser, M. Schlösser, L. Schlüter, H. Schön, K. Schönung, M. Schrank, B. Schulz, J. Schwarz, H. Seitz-Moskaliuk, W. Seller, V. Sibille, D. Siegmann, A. Skasyrskaya, M. Slezák, A. Špalek, F. Spanier, M. Steidl, N. Steinbrink, M. Sturm, M. Suesser, M. Sun, D. Tcherniakhovski, H. H. Telle, T. Thümmler, L. A. Thorne, N. Titov, I. Tkachev, N. Trost, K. Urban, D. Vénos, K. Valerius, B. A. Vandevender, R. Vianden, A. P. Vizcaya Hernández, B. L. Wall, S. Wüstling, M. Weber, C. Weinheimer, C. Weiss, S. Welte, J. Wendel, K. J. Wierman, J. F. Wilkerson, J. Wolf, W. Xu, Y. R. Yen, M. Zacher, S. Zadorozhny, M. Zbořil, and G. Zeller, Physical Review Letters 123, 1 (2019), arXiv:1909.06048 .
- Aker et al. (2022) M. Aker, A. Beglarian, J. Behrens, A. Berlev, U. Besserer, B. Bieringer, F. Block, B. Bornschein, L. Bornschein, M. Böttcher, T. Brunst, T. S. Caldwell, R. M. D. Carney, L. L. Cascio, S. Chilingaryan, W. Choi, K. Debowski, M. Deffert, M. Descher, D. D. Barrero, P. J. Doe, O. Dragoun, G. Drexlin, K. Eitel, E. Ellinger, R. Engel, S. Enomoto, A. Felden, J. A. Formaggio, F. M. Fränkle, G. B. Franklin, F. Friedel, A. Fulst, K. Gauda, W. Gil, F. Glück, R. Grössle, R. Gumbsheimer, V. Gupta, T. Höhn, V. Hannen, N. Haußmann, K. Helbing, S. Hickford, R. Hiller, D. Hillesheimer, D. Hinz, T. Houdy, A. Huber, A. Jansen, C. Karl, F. Kellerer, J. Kellerer, M. Klein, C. Köhler, L. Köllenberger, A. Kopmann, M. Korzeczek, A. Kovalík, B. Krasch, H. Krause, N. Kunka, T. Lasserre, T. L. Le, O. Lebeda, B. Lehnert, A. Lokhov, M. Machatschek, E. Malcherek, M. Mark, A. Marsteller, E. L. Martin, C. Melzer, A. Menshikov, S. Mertens, J. Mostafa, K. Müller, S. Niemes, P. Oelpmann, D. S. Parno, A. W. P. Poon, J. M. L. Poyato, F. Priester, M. Röllig, C. Röttele, R. G. H. Robertson, W. Rodejohann, C. Rodenbeck, M. Ryšavý, R. Sack, A. Saenz, P. Schäfer, A. Schaller, L. Schimpf, K. Schlösser, M. Schlösser, L. Schlüter, S. Schneidewind, M. Schrank, B. Schulz, A. Schwemmer, M. Šefĉík, V. Sibille, D. Siegmann, M. Slezák, M. Steidl, M. Sturm, M. Sun, D. Tcherniakhovski, H. H. Telle, L. A. Thorne, T. Thümmler, N. Titov, I. Tkachev, K. Urban, K. Valerius, D. Vénos, A. P. V. Hernández, C. Weinheimer, S. Welte, J. Wendel, J. F. Wilkerson, J. Wolf, S. Wüstling, W. Xu, Y. R. Yen, S. Zadoroghny, and G. Zeller, Nature Physics 18, 160 (2022).
- KATRIN Collaboration et al. (2025) KATRIN Collaboration et al., Science 388, 180 (2025).
- Esfahani et al. (2017) A. A. Esfahani, D. M. Asner, S. Böser, R. Cervantes, C. Claessens, L. de Viveiros, P. J. Doe, S. Doeleman, J. L. Fernandes, M. Fertl, E. C. Finn, J. A. Formaggio, D. Furse, M. Guigue, K. M. Heeger, A. M. Jones, K. Kazkaz, J. A. Kofron, C. Lamb, B. H. LaRoque, E. Machado, E. L. McBride, M. L. Miller, B. Monreal, P. Mohanmurthy, J. A. Nikkel, N. S. Oblath, W. C. Pettus, R. G. H. Robertson, L. J. Rosenberg, G. Rybka, D. Rysewyk, L. Saldaña, P. L. Slocum, M. G. Sternberg, J. R. Tedeschi, T. Thümmler, B. A. VanDevender, L. E. Vertatschitsch, M. Wachtendonk, J. Weintroub, N. L. Woods, A. Young, and E. M. Zayas, Journal of Physics G: Nuclear and Particle Physics 44, 054004 (2017).
- Betti et al. (2019) M. Betti, M. Biasotti, A. Boscá, F. Calle, N. Canci, G. Cavoto, C. Chang, A. Cocco, A. Colijn, J. Conrad, N. D’Ambrosio, N. D. Groot, P. de Salas, M. Faverzani, A. Ferella, E. Ferri, P. Garcia-Abia, I. García-Cortés, G. G. Gomez-Tejedor, S. Gariazzo, F. Gatti, C. Gentile, A. Giachero, J. Gudmundsson, Y. Hochberg, Y. Kahn, A. Kievsky, M. Lisanti, C. Mancini-Terracciano, G. Mangano, L. Marcucci, C. Mariani, J. Martínez, M. Messina, A. Molinero-Vela, E. Monticone, A. Moroño, A. Nucciotti, F. Pandolfi, S. Parlati, S. Pastor, J. Pedrós, C. P. de los Heros, O. Pisanti, A. Polosa, A. Puiu, I. Rago, Y. Raitses, M. Rajteri, N. Rossi, I. Rucandio, R. Santorelli, K. Schaeffner, C. Tully, M. Viviani, F. Zhao, and K. Zurek, Journal of Cosmology and Astroparticle Physics 2019, 047 (2019).
- Gastaldo et al. (2014) L. Gastaldo, K. Blaum, A. Doerr, C. E. Düllmann, K. Eberhardt, S. Eliseev, C. Enss, A. Faessler, A. Fleischmann, S. Kempf, M. Krivoruchenko, S. Lahiri, M. Maiti, Y. N. Novikov, P. C. Ranitzsch, F. Simkovic, Z. Szusc, and M. Wegner, Journal of Low Temperature Physics 176, 876 (2014), arXiv:1306.2655 .
- Gastaldo et al. (2017) L. Gastaldo, K. Blaum, K. Chrysalidis, T. Day Goodacre, A. Domula, M. Door, H. Dorrer, C. E. Düllmann, K. Eberhardt, S. Eliseev, C. Enss, A. Faessler, P. Filianin, A. Fleischmann, D. Fonnesu, L. Gamer, R. Haas, C. Hassel, D. Hengstler, J. Jochum, K. Johnston, U. Kebschull, S. Kempf, T. Kieck, U. Köster, S. Lahiri, M. Maiti, F. Mantegazzini, B. Marsh, P. Neroutsos, Y. N. Novikov, P. C. Ranitzsch, S. Rothe, A. Rischka, A. Saenz, O. Sander, F. Schneider, S. Scholl, R. X. Schüssler, C. Schweiger, F. Simkovic, T. Stora, Z. Szücs, A. Türler, M. Veinhard, M. Weber, M. Wegner, K. Wendt, and K. Zuber, European Physical Journal: Special Topics 226, 1623 (2017).
- Velte et al. (2019) C. Velte, F. Ahrens, A. Barth, K. Blaum, M. Braß, M. Door, H. Dorrer, C. E. Düllmann, S. Eliseev, C. Enss, P. Filianin, A. Fleischmann, L. Gastaldo, A. Goeggelmann, T. D. Goodacre, M. W. Haverkort, D. Hengstler, J. Jochum, K. Johnston, M. Keller, S. Kempf, T. Kieck, C. M. König, U. Köster, K. Kromer, F. Mantegazzini, B. Marsh, Y. N. Novikov, F. Piquemal, C. Riccio, D. Richter, A. Rischka, S. Rothe, R. X. Schüssler, C. Schweiger, T. Stora, M. Wegner, K. Wendt, M. Zampaolo, and K. Zuber, The European Physical Journal C 79 (2019), 10.1140/epjc/s10052-019-7513-x.
- Cattadori et al. (2007) C. M. Cattadori, M. De Deo, M. Laubenstein, L. Pandola, and V. I. Tretyak, Physics of Atomic Nuclei 70, 127 (2007).
- Haaranen and Suhonen (2013) M. Haaranen and J. Suhonen, The European Physical Journal A 49, 1 (2013).
- Suhonen (2014) J. Suhonen, Physica Scripta 89, 54032 (2014).
- Ejiri et al. (2019) H. Ejiri, J. Suhonen, and K. Zuber, Physics Reports 797, 1 (2019).
- Sandler et al. (2019) R. Sandler, G. Bollen, N. D. Gamage, A. Hamaker, C. Izzo, D. Puentes, M. Redshaw, R. Ringle, and I. Yandow, Physical Review C 100, 1 (2019), arXiv:1906.03335 .
- Karthein et al. (2019) J. Karthein, D. Atanasov, K. Blaum, S. Eliseev, P. Filianin, D. Lunney, V. Manea, M. Mougeot, D. Neidherr, Y. Novikov, L. Schweikhard, A. Welker, F. Wienholtz, and K. Zuber, Hyperfine Interactions 240, 1 (2019), arXiv:1905.05510 .
- De Roubin et al. (2020) A. De Roubin, J. Kostensalo, T. Eronen, L. Canete, R. P. De Groote, A. Jokinen, A. Kankainen, D. A. Nesterenko, I. D. Moore, S. Rinta-Antila, J. Suhonen, and M. Vilén, Physical Review Letters 124, 1 (2020), arXiv:2002.08282 .
- Ge et al. (2021a) Z. Ge, T. Eronen, A. de Roubin, D. A. Nesterenko, M. Hukkanen, O. Beliuskina, R. de Groote, S. Geldhof, W. Gins, A. Kankainen, A. Koszorús, J. Kotila, J. Kostensalo, I. D. Moore, A. Raggio, S. Rinta-Antila, J. Suhonen, V. Virtanen, A. P. Weaver, A. Zadvornaya, and A. Jokinen, Physical Review C 103, 065502 (2021a).
- Ge et al. (2021b) Z. Ge, T. Eronen, K. S. Tyrin, J. Kotila, J. Kostensalo, D. A. Nesterenko, O. Beliuskina, R. de Groote, A. de Roubin, S. Geldhof, W. Gins, M. Hukkanen, A. Jokinen, A. Kankainen, A. Koszorús, M. I. Krivoruchenko, S. Kujanpää, I. D. Moore, A. Raggio, S. Rinta-Antila, J. Suhonen, V. Virtanen, A. P. Weaver, and A. Zadvornaya, Phys. Rev. Lett. 127, 272301 (2021b).
- Ge et al. (2022a) Z. Ge, T. Eronen, A. de Roubin, K. Tyrin, L. Canete, S. Geldhof, A. Jokinen, A. Kankainen, J. Kostensalo, J. Kotila, M. Krivoruchenko, I. Moore, D. Nesterenko, J. Suhonen, and M. Vilén, Physics Letters B 832, 137226 (2022a).
- Eronen et al. (2022) T. Eronen, Z. Ge, A. de Roubin, M. Ramalho, J. Kostensalo, J. Kotila, O. Beliushkina, C. Delafosse, S. Geldhof, W. Gins, M. Hukkanen, A. Jokinen, A. Kankainen, I. Moore, D. Nesterenko, M. Stryjczyk, and J. Suhonen, Physics Letters B 830, 137135 (2022).
- Ge et al. (2022b) Z. Ge, T. Eronen, A. de Roubin, J. Kostensalo, J. Suhonen, D. A. Nesterenko, O. Beliuskina, R. de Groote, C. Delafosse, S. Geldhof, W. Gins, M. Hukkanen, A. Jokinen, A. Kankainen, J. Kotila, A. Koszorús, I. D. Moore, A. Raggio, S. Rinta-Antila, V. Virtanen, A. P. Weaver, and A. Zadvornaya, Phys. Rev. C 106, 015502 (2022b).
- Ramalho et al. (2022) M. Ramalho, Z. Ge, T. Eronen, D. A. Nesterenko, J. Jaatinen, A. Jokinen, A. Kankainen, J. Kostensalo, J. Kotila, M. I. Krivoruchenko, J. Suhonen, K. S. Tyrin, and V. Virtanen, Phys. Rev. C 106, 015501 (2022).
- Gamage et al. (2022) N. D. Gamage, R. Sandler, F. Buchinger, J. A. Clark, D. Ray, R. Orford, W. S. Porter, M. Redshaw, G. Savard, K. S. Sharma, and A. A. Valverde, Phys. Rev. C 106, 045503 (2022).
- Keblbeck et al. (2023) D. K. Keblbeck, R. Bhandari, N. D. Gamage, M. Horana Gamage, K. G. Leach, X. Mougeot, and M. Redshaw, Phys. Rev. C 107, 015504 (2023).
- Ge et al. (2023) Z. Ge, T. Eronen, A. de Roubin, M. Ramalho, J. Kostensalo, J. Kotila, J. Suhonen, D. A. Nesterenko, A. Kankainen, P. Ascher, O. Beliuskina, M. Flayol, M. Gerbaux, S. Grévy, M. Hukkanen, A. Husson, A. Jaries, A. Jokinen, I. D. Moore, P. Pirinen, J. Romero, M. Stryjczyk, V. Virtanen, and A. Zadvornaya, Phys. Rev. C 108, 045502 (2023).
- Valverde et al. (2024) A. Valverde, F. Kondev, B. Liu, D. Ray, M. Brodeur, D. Burdette, N. Callahan, A. Cannon, J. Clark, D. Hoff, R. Orford, W. Porter, G. Savard, K. Sharma, and L. Varriano, Physics Letters B 858, 139037 (2024).
- Ge et al. (2024a) Z. Ge, T. Eronen, M. Ramalho, A. de Roubin, D. A. Nesterenko, A. Kankainen, O. Beliuskina, R. de Groote, S. Geldhof, W. Gins, M. Hukkanen, A. Jokinen, Á. Koszorús, J. Kotila, J. Kostensalo, I. D. Moore, P. Pirinen, A. Raggio, S. Rinta-Antila, V. A. Sevestrean, J. Suhonen, V. Virtanen, and A. Zadvornaya, Eur. Phys. J. A 60 (2024a), https://doi.org/10.1140/epja/s10050-024-01317-3.
- Ge et al. (2024b) Z. Ge, T. Eronen, V. A. Sevestrean, O. Niţescu, S. Stoica, M. Ramalho, J. Suhonen, A. de Roubin, D. Nesterenko, A. Kankainen, P. Ascher, S. Ayet San Andres, O. Beliuskina, P. Delahaye, M. Flayol, M. Gerbaux, S. Grévy, M. Hukkanen, A. Jaries, A. Jokinen, A. Husson, D. Kahl, J. Kostensalo, J. Kotila, I. Moore, S. Nikas, M. Stryjczyk, and V. Virtanen, Physics Letters B 859, 139094 (2024b).
- NND (2021) “National nuclear data center,” Available at https://www.nndc.bnl.gov/ (2020/4/7) (2021).
- Wang et al. (2021) M. Wang, W. Huang, F. Kondev, G. Audi, and S. Naimi, Chinese Physics C 45, 030003 (2021).
- Moore et al. (2013) I. D. Moore, T. Eronen, D. Gorelov, J. Hakala, A. Jokinen, A. Kankainen, V. S. Kolhinen, J. Koponen, H. Penttilä, I. Pohjalainen, M. Reponen, J. Rissanen, A. Saastamoinen, S. Rinta-Antila, V. Sonnenschein, and J. Äystö, Nuclear Instruments and Methods in Physics Research, Section B: Beam Interactions with Materials and Atoms 317, 208 (2013).
- Kolhinen et al. (2013) V. S. Kolhinen, T. Eronen, D. Gorelov, J. Hakala, A. Jokinen, K. Jokiranta, A. Kankainen, M. Koikkalainen, J. Koponen, H. Kulmala, M. Lantz, A. Mattera, I. D. Moore, H. Penttilä, T. Pikkarainen, I. Pohjlainen, M. Reponen, S. Rinta-Antila, J. Rissanen, C. Rodríguez Triguero, K. Rytkönen, A. Saastamoinen, A. Solders, V. Sonnenschein, and J. Äystö, Nuclear Instruments and Methods in Physics Research, Section B: Beam Interactions with Materials and Atoms 317, 506 (2013).
- Karvonen et al. (2008) P. Karvonen, I. D. Moore, T. Sonoda, T. Kessler, H. Penttilä, K. Peräjärvi, P. Ronkanen, and J. Äystö, Nuclear Instruments and Methods in Physics Research, Section B: Beam Interactions with Materials and Atoms 266, 4794 (2008).
- Nieminen et al. (2001) A. Nieminen, J. Huikari, A. Jokinen, J. Äystö, P. Campbell, and E. C. Cochrane, Nuclear Instruments and Methods in Physics Research, Section A: Accelerators, Spectrometers, Detectors and Associated Equipment 469, 244 (2001).
- Vilén et al. (2020) M. Vilén, L. Canete, B. Cheal, A. Giatzoglou, R. de Groote, A. de Roubin, T. Eronen, S. Geldhof, A. Jokinen, A. Kankainen, I. D. Moore, D. A. Nesterenko, H. Penttilä, I. Pohjalainen, M. Reponen, and S. Rinta-Antila, Nuclear Instruments and Methods in Physics Research, Section B: Beam Interactions with Materials and Atoms 463, 382 (2020).
- Eronen et al. (2012) T. Eronen, V. S. Kolhinen, V. V. Elomaa, D. Gorelov, U. Hager, J. Hakala, A. Jokinen, A. Kankainen, P. Karvonen, S. Kopecky, I. D. Moore, H. Penttilä, S. Rahaman, S. Rinta-Antila, J. Rissanen, A. Saastamoinen, J. Szerypo, C. Weber, and J. Äystö, European Physical Journal A 48, 1 (2012).
- Savard et al. (1991) G. Savard, S. Becker, G. Bollen, H. J. Kluge, R. B. Moore, T. Otto, L. Schweikhard, H. Stolzenberg, and U. Wiess, Physics Letters A 158, 247 (1991).
- Eronen et al. (2008) T. Eronen, V.-V. Elomaa, U. Hager, J. Hakala, A. Jokinen, A. Kankainen, S. Rahaman, J. Rissanen, C. Weber, and J. Äystö, Nuclear Instruments and Methods in Physics Research Section B: Beam Interactions with Materials and Atoms 266, 4527 (2008), proceedings of the XVth International Conference on Electromagnetic Isotope Separators and Techniques Related to their Applications.
- Nesterenko et al. (2021) D. A. Nesterenko, T. Eronen, Z. Ge, A. Kankainen, and M. Vilen, Eur. Phys. J. A 57, 302 (2021).
- Nesterenko et al. (2018) D. A. Nesterenko, T. Eronen, A. Kankainen, L. Canete, A. Jokinen, I. D. Moore, H. Penttilä, S. Rinta-Antila, A. de Roubin, and M. Vilen, European Physical Journal A 54, 0 (2018).
- Eliseev et al. (2014) S. Eliseev, K. Blaum, M. Block, A. Dörr, C. Droese, T. Eronen, M. Goncharov, M. Höcker, J. Ketter, E. M. Ramirez, D. A. Nesterenko, Y. N. Novikov, and L. Schweikhard, Applied Physics B: Lasers and Optics 114, 107 (2014).
- Eliseev et al. (2013) S. Eliseev, K. Blaum, M. Block, C. Droese, M. Goncharov, E. Minaya Ramirez, D. A. Nesterenko, Y. N. Novikov, and L. Schweikhard, Physical Review Letters 110, 82501 (2013).
- Kramida et al. (2020) A. Kramida, Yu. Ralchenko, J. Reader, and and NIST ASD Team, NIST Atomic Spectra Database (ver. 5.8), [Online]. Available: https://physics.nist.gov/asd [2021, January 19]. National Institute of Standards and Technology, Gaithersburg, MD. (2020).
- X-R (2021) “X-ray data booklet,” Available at https://xdb.lbl.gov (2021/05/01) (2021).
- Kellerbauer et al. (2003) A. Kellerbauer, K. Blaum, G. Bollen, F. Herfurth, H. J. Kluge, M. Kuckein, E. Sauvan, C. Scheidenberger, and L. Schweikhard, European Physical Journal D 22, 53 (2003).
- Roux et al. (2013) C. Roux, K. Blaum, M. Block, C. Droese, S. Eliseev, M. Goncharov, F. Herfurth, E. M. Ramirez, D. A. Nesterenko, Y. N. Novikov, and L. Schweikhard, The European Physical Journal D 67, 1 (2013).
- Birge (1932) R. T. Birge, Physical Review 40, 207 (1932).
- Huang et al. (2021) W. Huang, M. Wang, F. Kondev, G. Audi, and S. Naimi, Chinese Physics C 45, 030002 (2021).
- Salvat and Fernández-Varea (2019) F. Salvat and J. M. Fernández-Varea, Computer Physics Communications 240, 165 (2019).
- Brown and Rae (2014) B. Brown and W. Rae, Nuclear Data Sheets 120, 115 (2014).
- Mougeot (2018) X. Mougeot, Applied Radiation and Isotopes 134, 225 (2018).
- Sevestrean et al. (2023) V. A. Sevestrean, O. Niţescu, S. Ghinescu, and S. Stoica, Phys. Rev. A 108, 012810 (2023).
- CAMPBELL and PAPP (2001) J. CAMPBELL and T. PAPP, Atomic Data and Nuclear Data Tables 77, 1 (2001).
- Bambynek et al. (1977) W. Bambynek, H. Behrens, M. H. Chen, B. Crasemann, M. L. Fitzpatrick, K. W. D. Ledingham, H. Genz, M. Mutterer, and R. L. Intemann, Rev. Mod. Phys. 49, 77 (1977).
- Koning, Arjan et al. (2023) Koning, Arjan, Hilaire, Stephane, and Goriely, Stephane, Eur. Phys. J. A 59, 131 (2023).
- Schweiger et al. (2024) C. Schweiger, M. Braß, V. Debierre, M. Door, H. Dorrer, C. E. Düllmann, C. Enss, P. Filianin, L. Gastaldo, Z. Harman, M. W. Haverkort, J. Herkenhoff, P. Indelicato, C. H. Keitel, K. Kromer, D. Lange, Y. N. Novikov, D. Renisch, A. Rischka, R. X. Schüssler, S. Eliseev, and K. Blaum, Nature Physics 20, 921 (2024).
- Gastaldo (2022) L. Gastaldo, Journal of Low Temperature Physics 209, 804 (2022).